OpenStudy (ethan01):

Solve the following system of equations. 2x + y = 3 x = 2y - 1

OpenStudy (yanasidlinskiy):

Plug it in: 2(2y-1)+y = 3 Solve for y Plug that in for the second equation. Right @ganeshie8 ??

ganeshie8 (ganeshie8):

Looks perfect !

OpenStudy (ethan01):

thanks can you help me with one more

OpenStudy (yanasidlinskiy):

Yep:)

OpenStudy (ethan01):

Solve the following system of equations. 3x + 2y - 5 = 0 x = y + 10

OpenStudy (yanasidlinskiy):

Sorry Ethan. Almost reported you hahha by accident lol. Same thing. Just plug it in and solve ;) 3(y+10)+2y-5=0 Right?

OpenStudy (ethan01):

thanks

OpenStudy (yanasidlinskiy):

Yeap anytime ;)